Project

General

Profile

NDOS Commissioning Meeting 2012/01/26

Attendance: Jon, Mark Mathis, Jan, Nick, Susan, Leon, Mat, Gavin, Athans,

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the next week
  3. DAQ Update and Plans for next week
  4. DCS Update and Plans for next week
  5. AOB

Beam and Detector Report

  • Jon's Report: unable to make plots from nearline due to dB downtime.
  • In general the detector showed no significant issues this past week.

Installation plans for next week

  • Hopefully next week there will some tests on the APD purge system. Mat
  • Jon: what is that status of the APDs from Hamamatsu? Leon: were supposed to have shipped today, but haven't heard confirmation of this so probably not until next week.
  • Mat: how many will be shipped. Leon: probably 50-100 units.
  • Jon: how long for turnaround? Leon: about a week.

DAQ Update and Plans for next week

  • RC updates... problem with dB writes, so there are missing rows for runs between this past Tuesday and today.
  • DB has been migrated...
  • Plans for tomorrow...

DCS Update and Plans for next week

AOB

  • New meeting time? UMN will not be able to join us starting next week, and there are others who are generally unable to attend at this time. Monday afternoons might be a good